OpenStudy (anonymous):

1 3/4 divided by 3 = ?

OpenStudy (anonymous):

21/4 3 1/4 1 4/7 7/12

OpenStudy (alexandervonhumboldt2):

first convert into improper fraction:\ \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[1\frac{ 3 }{ 4 }=\frac{ 1*4+3 }{ 4 }=\frac{ 7 }{ 4 }\] now divide: \[\frac{ 7 }{ 4 }/3=\frac{ 7 }{ 4 }*\frac{ 1 }{ 3 }=\frac{ 7 }{ 12 }\]